0.7n-15+n=2n-8-0.4n and can you please show me how to check it, thanks
Answer by edjones(8007)

You have to solve it before you can check it.
0.7n-15+n=2n-8-0.4n
combine terms: 1.7n-15=1.6n-8
Subtract 1.6n from both sides. 0.1n-15=-8
add 15 to both sides. 0.1n=7
divide 0.1 into both sides. n=70 Ans.
Now we check by replacing n every place we find it in the original equation with 70.
0.7*70-15+70=2*70-8-.4*70
49-15+70=140-8-28
104=104
EdJones